### v1

# MOR-VIT: Efficient Vision Transformer with Mixture-of-Recursions

Abstract
Vision Transformers (ViTs) have achieved remarkable success in image recognition, yet standard ViT architectures are hampered by substantial parameter redundancy and high computational cost, limiting their practical deployment. While recent efforts on efficient ViTs primarily focus on static model compression or token-level sparsification, they remain constrained by fixed computational depth for all tokens. In this work, we present MoR-ViT, a novel vision transformer framework that, for the first time, incorporates a token-level dynamic recursion mechanism inspired by the Mixture-of-Recursions (MoR) paradigm. This approach enables each token to adaptively determine its processing depth, yielding a flexible and input-dependent allocation of computational resources. Extensive experiments on ImageNet-1K and transfer benchmarks demonstrate that MoR-ViT not only achieves state-of-the-art accuracy with up to 70% parameter reduction and 2.5× inference acceleration, but also outperforms leading efficient ViT baselines such as DynamicViT and TinyViT under comparable conditions. These results establish dynamic recursion as an effective strategy for efficient vision transformers and open new avenues for scalable and deployable deep learning models in real-world scenarios.
1 Introduction
Vision Transformers (ViTs) have emerged as a powerful paradigm in visual recognition, rivaling and in some cases surpassing traditional convolutional neural networks (CNNs) such as ResNet [1]. The seminal Vision Transformer (ViT) [2] models images as sequences of patch tokens and applies the standard Transformer encoder, achieving strong performance across various benchmarks. However, vanilla ViTs are characterized by substantial parameter redundancy and high computational costs, limiting their efficiency and real-world deployment.
To address these limitations, the field has witnessed a surge of research on efficient Vision Transformer architectures in recent years. DynamicViT [3] introduced token-level dynamic sparsification, adaptively dropping uninformative tokens at each layer to reduce computation. Methods such as Adaptive Token Sampling (ATS) [4], EViT [5], and related dynamic token pruning techniques further refine this direction, enabling input-adaptive computation and maintaining high accuracy at reduced FLOPs. EfficientFormer [6] and PoolFormer (MetaFormer) [7] optimize network structures and operators for latency and parameter efficiency, demonstrating that Transformer-based models can achieve mobile-level speed and compactness. MobileViT [8] combines convolutional and Transformer modules to deliver lightweight, mobile-friendly backbones. Recent works on TinyViT [9] leverage large-scale distillation and optimized architectures to push the accuracy of small ViTs, rivaling or outperforming much larger models in compact settings. Other innovations, such as TokenLearner [10], propose dynamic token aggregation modules that learn to summarize global information with a small set of latent tokens, while SparseFormer [11] explores extremely sparse visual representations using limited learnable queries.
Despite significant advances, nearly all existing efficient ViT methods operate under a key constraint: they apply fixed computational depth and identical processing to all tokens, regardless of semantic complexity or importance. As a result, even highly-optimized ViTs may underutilize potential savings by failing to allocate computation in a token-adaptive, recursive manner.
To address this gap, we propose MoR-ViT, a novel Vision Transformer architecture that, for the first time, incorporates a token-level dynamic recursion mechanism inspired by the Mixture-of-Recursions (MoR) paradigm. In MoR-ViT, each token is routed through the network with a flexible, input-dependent recursive depth determined by a lightweight router, enabling adaptive allocation of computation based on token semantics. Our experiments on ImageNet-1K and transfer learning benchmarks show that MoR-ViT achieves competitive or superior accuracy to state-of-the-art efficient ViTs—including DynamicViT, TinyViT, and others—while delivering up to 70% parameter reduction and 2.5$\times$ inference acceleration. These results highlight dynamic recursion as a new and effective strategy for scalable, resource-efficient visual modeling, opening promising avenues for future research and deployment.
Notably, MoR-ViT achieves these improvements without relying on additional domain-specific resources such as large-scale distillation or massive external pretraining data, and in several cases, even surpasses baseline performance on key evaluation metrics.
2 Related Work
2.1 Comprehensive Surveys on Vision Transformers
A number of comprehensive surveys have recently reviewed the rapid evolution of vision transformers (ViTs) and their extensions [12, 13]. These works systematically categorize visual transformer architectures, analyze self-attention variants, hybrid structures, and summarize applications in classification, detection, and generative modeling. They also discuss open challenges in efficiency, scalability, and deployment, highlighting the need for further research in resource-constrained and adaptive computation settings.
2.2 Token Sparsification and Dynamic Token Pruning Methods
A prominent line of research focuses on reducing redundancy in ViTs by adaptively selecting or pruning input tokens. DynamicViT [3] introduced a dynamic token sparsification mechanism, allowing the model to drop uninformative tokens at each layer. Adaptive Token Sampling (ATS) [4] and EViT [5] further refine token pruning by learning to select salient patches based on content. TokenLearner [10] and SparseFormer [11] explore token aggregation and sparse latent token representations, enabling highly efficient global information modeling. These approaches have achieved significant computation savings and maintained high accuracy, but still apply uniform processing depth to all tokens.
2.3 Efficient Architecture Design and Model Compression
Another major direction is the design of efficient transformer architectures and parameter compression. EfficientFormer [6] and PoolFormer (MetaFormer) [7] propose operator- and structure-level optimizations, drastically reducing inference latency and model size. MobileViT [8] integrates convolutional priors with transformer modules for mobile-friendly vision backbones. TinyViT [9] employs large-scale distillation and compact design to deliver state-of-the-art accuracy with minimal parameters. While these works demonstrate that ViTs can be made lightweight and fast, they still process all tokens with fixed computational schedules.
2.4 Limitations of Existing Methods and Motivation for This Work
Despite remarkable progress, nearly all existing efficient ViT approaches share a key limitation: they apply fixed computational depth and identical processing to all tokens, regardless of their individual semantic importance or complexity. This uniformity limits resource allocation flexibility and may result in unnecessary computation on trivial regions. Our work addresses this gap by introducing MoR-ViT, which enables token-wise dynamic recursion and adaptive depth allocation within the transformer, leading to more resource-efficient visual modeling.
3 Method
3.1 Vision Transformer (ViT) Overview
The architecture and main computational flow of the Vision Transformer (ViT) [2] are shown in Figure 1. ViT applies the transformer framework directly to image recognition.
Given an input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, we split it into $N$ non-overlapping patches, each of size $P\times P$. Each patch $\mathbf{I}_{p}$ is flattened into a vector and mapped to a $D$-dimensional token embedding via a learnable linear projection $\mathbf{E}\in\mathbb{R}^{(P^{2}\cdot 3)\times D}$:
$\mathbf{x}_{p}=\text{Flatten}(\mathbf{I}_{p})\cdot\mathbf{E}\in\mathbb{R}^{D},\quad p=1,\ldots,N$
The sequence of patch embeddings is denoted as $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{N}]\in\mathbb{R}^{N\times D}$. A class token $\mathbf{x}_{\text{cls}}\in\mathbb{R}^{D}$ is prepended, and learnable positional embeddings $\mathbf{E}_{\text{pos}}\in\mathbb{R}^{(N+1)\times D}$ are added:
$\mathbf{Z}_{0}=[\mathbf{x}_{\text{cls}};\mathbf{x}_{1};\ldots;\mathbf{x}_{N}]+\mathbf{E}_{\text{pos}}$
The resulting sequence $\mathbf{Z}_{0}$ is then fed into a stack of $L$ standard Transformer encoder blocks.
Figure 1: Architecture of the Vision Transformer (ViT). The input image is split into patches, linearly embedded, concatenated with a class token, and augmented with positional encodings. The resulting sequence is processed by transformer encoder layers, and the class token output is used for final classification.
3.2 Mixture-of-Recursions (MoR) Mechanism
The Mixture-of-Recursions (MoR) framework builds upon the standard Transformer pipeline by incorporating a dynamic routing mechanism that adaptively allocates recursive computation for each token. The input embedding, patch partitioning, and positional encoding are identical to the vanilla ViT and Transformer architectures.
Given an input token $\mathbf{x}_{t}$ (e.g., a patch embedding with position encoding), MoR employs a lightweight router to determine the recursion depth assigned to each token. At each recursion step $r$, the router computes a routing score for token $t$:
$g^{r}_{t}=\mathcal{G}(\theta_{r}^{\top}\mathbf{h}^{r}_{t})$
where $\mathbf{h}^{r}_{t}$ is the hidden state of token $t$ at recursion step $r$, $\theta_{r}$ is the router parameter, and $\mathcal{G}(\cdot)$ is an activation function (e.g., sigmoid or tanh).
A $\beta$-percentile threshold $P_{\beta}(G^{r})$ is computed over all routing scores at step $r$, and tokens with $g^{r}_{t}>P_{\beta}(G^{r})$ are selected to continue to the next recursion step, while others exit:
$\mathbf{h}^{r+1}_{t}=\begin{cases}g^{r}_{t}f(\mathbf{h}^{r}_{t},\Phi^{\prime})+\mathbf{h}^{r}_{t},&\text{if }g^{r}_{t}>P_{\beta}(G^{r})\\ \mathbf{h}^{r}_{t},&\text{otherwise}\end{cases}$
Here, $f(\cdot)$ is the shared Transformer block, and $\Phi^{\prime}$ are its parameters. The mask matrix for token selection is updated at each recursion step. For implementation efficiency, only the selected tokens’ key-value (KV) pairs are cached and participate in subsequent self-attention computations, following the practice in [14].
[…]
The objective function consists of the main task loss (e.g., cross-entropy for classification) and possible auxiliary losses to encourage stable routing and balanced computation.
Notation Summary:
  • $g^{r}_{t}$: routing score for token $t$ at recursion $r$
  • $P_{\beta}(G^{r})$: $\beta$-percentile threshold for routing at step $r$
  • $\mathbf{h}^{r}_{t}$: hidden state at recursion $r$
  • $\Phi^{\prime}$: shared parameters of recursion block
  • $f(\cdot)$: transformer block function
Figure 2 illustrates the overall workflow of the Mixture-of-Recursions (MoR) mechanism, which enables token-wise adaptive computation by dynamically assigning different recursion depths to each input token.
Figure 2: Overview of the Mixture-of-Recursions (MoR) mechanism. Each token is assigned a dynamic recursion depth by the router, leading to token-wise adaptive computation. The router computes gating scores for each token, constructs a dynamic binary mask based on the $\beta$-th percentile, and enables token-specific exit from the recursion block.
3.3 MoR-ViT Architecture
The overall workflow of the proposed MoR-ViT architecture, which integrates ViT-style patch embedding with the Mixture-of-Recursions (MoR) mechanism to enable token-wise adaptive computation for image recognition, is illustrated in Figure 3.
Given an input image $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$, the image is first divided into $N$ non-overlapping patches of size $P\times P$, where $N=HW/P^{2}$. Each patch is flattened and projected into a $D$-dimensional token embedding:
$\mathbf{x}_{p}=\text{Flatten}(\mathbf{X}_{p})\,\mathbf{W}_{e}+\mathbf{b}_{e},\quad p=1,\ldots,N$
A learnable class token $\mathbf{x}_{\mathrm{cls}}\in\mathbb{R}^{D}$ is prepended, and positional embeddings $\mathbf{E}_{\mathrm{pos}}\in\mathbb{R}^{(N+1)\times D}$ are added:
$\mathbf{Z}_{0}=[\mathbf{x}_{\mathrm{cls}};\mathbf{x}_{1};\ldots;\mathbf{x}_{N}]+\mathbf{E}_{\mathrm{pos}}$
The sequence $\mathbf{Z}_{0}$ serves as the input tokens for the MoR module.
Within the MoR block, each token undergoes a shared transformation. At each recursion step $l$, a lightweight router computes a gating score for each token:
$g_{t}^{(l)}=\sigma\left(\mathbf{w}^{\top}h_{t}^{(l)}+b\right)$
where $h_{t}^{(l)}$ is the hidden state of token $t$ at step $l$, and $\sigma(\cdot)$ denotes the sigmoid activation.
To dynamically select tokens for further recursion, a binary mask is constructed by thresholding the gating scores at the adaptive $\beta$-th percentile:
$m_{t}^{(l)}=\begin{cases}0,&\text{if }g_{t}^{(l)}<\mathrm{Percentile}_{\beta}(\{g_{t^{\prime}}^{(l)}\}_{t^{\prime}=1}^{N})\\ 1,&\text{otherwise}\end{cases}$
Tokens with $m_{t}^{(l)}=0$ exit recursion at step $l$ and their outputs are preserved, while tokens with $m_{t}^{(l)}=1$ are recursively processed in the next step. This dynamic process allocates deeper recursion to more ambiguous tokens and allows simpler tokens to exit early.
For implementation efficiency, only the selected tokens’ key-value pairs are cached for subsequent self-attention computations, following [14].
After all recursion steps, the outputs from all tokens are aggregated, and the final [CLS] token is passed through a linear classifier (e.g., MLP) to produce class logits. The training objective is defined as:
$\mathcal{L}=\mathcal{L}_{\mathrm{task}}+\lambda\mathcal{L}_{\mathrm{routing}}$
where $\mathcal{L}_{\mathrm{task}}$ is the standard task loss (e.g., cross-entropy), $\mathcal{L}_{\mathrm{routing}}$ is a routing regularization term, and $\lambda$ is the trade-off coefficient. For image classification tasks such as CIFAR-10/100 and Oxford Flowers-102, the protocol follows ViT/DeiT [2, 15].
Figure 3: Architecture of MoR-ViT. The input image is split into patches, embedded as tokens, and processed through the MoR block, where each token is adaptively assigned a recursion depth by the router and aggregated for final classification.
Interpretability of Dynamic Routing.
To provide a deeper understanding of how our dynamic routing mechanism allocates computational depth at the token level, we visualize and contrast two routing scenarios: (1) the typical case where dynamic routing is effective, and (2) a degenerate case where all tokens are assigned uniformly shallow recursion depths.
Figure 4 presents the dynamic routing scenario. The left subfigure displays the spatial map of token recursion depths for a representative image, where warm colors (e.g., red) indicate higher recursion depth and cool colors (e.g., blue) indicate lower depth. The right subfigure shows the corresponding histogram, clearly demonstrating that most tokens are processed with shallow recursion, while a significant fraction are dynamically allocated to deeper computation. Notably, tokens in semantically complex or boundary regions are more likely to receive higher recursion depth, confirming the ability of MoR-ViT to perform adaptive, fine-grained computation according to token difficulty.
In contrast, Figure 5 depicts a degenerate scenario where almost all tokens are assigned only minimal recursion depth, and the dynamic mechanism effectively reduces to a uniform, static computation schedule. Both the spatial map and histogram in this case reveal an absence of adaptivity: nearly all tokens exit after the first recursion step, regardless of their semantic content.
This comparison highlights a key advantage of our approach: when the router successfully learns to differentiate token complexity, MoR-ViT achieves efficient resource allocation and strong generalization; when the router degenerates, potential efficiency gains are lost, resembling fixed-depth models. These visualizations provide compelling and intuitive evidence for the mechanism’s effectiveness, as well as a diagnostic tool for identifying failure cases or training pathologies in future research.
Figure 4: Dynamic routing example: (Left) Token-wise recursion depth map. Warm colors (red) indicate tokens assigned deeper recursion, cool colors (blue) indicate shallow recursion. (Right) Histogram confirms a clear dynamic allocation pattern, with a significant portion of tokens undergoing multiple recursion steps.
Figure 5: Degenerate routing example: (Left) Nearly all tokens are assigned shallow recursion (blue). (Right) Histogram reveals a uniform, non-adaptive allocation, demonstrating the mechanism’s limitation when the router fails to differentiate token difficulty.
4 Experiments
We systematically evaluate the representation capabilities and efficiency of MoRViT in comparison with a broad range of recent vision transformer architectures and efficient baselines. In addition to canonical models such as ResNet [1], Vision Transformer (ViT) [2], and Data-efficient Image Transformer (DeiT) [15], our evaluation includes comprehensive horizontal comparisons with leading efficient ViT variants, namely DynamicViT [3], Adaptive Token Sampling (ATS) [4], EfficientFormer [6], PoolFormer (MetaFormer) [7], MobileViT [8], TinyViT [9], TokenLearner [10], and SparseFormer [11]. These models represent state-of-the-art approaches in token sparsification, dynamic routing, architectural compression, and mobile-friendly transformer design.
To ensure fair and rigorous comparison, all transformer models are trained under standard supervised scenarios, exclusively on the ImageNet dataset [16], without the use of extra large-scale data or self-supervised pretraining. For downstream transfer learning evaluation, we follow established protocols [2, 15] and conduct fine-tuning as well as testing on public CIFAR [17] and Oxford Flowers [18] datasets. By benchmarking across this diverse set of models and benchmarks, we aim to highlight the advantages and trade-offs of MoRViT relative to both conventional and the latest efficient ViT-based methods.
4.1 Experiment Setup
Our model is built upon the standard ViT backbone (patch embedding + transformer encoder), with the Mixture-of-Recursions (MoR) mechanism seamlessly integrated to enhance the model’s adaptive capacity. This modular design preserves the advantages of ViT while introducing substantial novelty through the MoR module. All models are trained from scratch on the ImageNet dataset [16] using supervised learning, without any external large-scale data or self-supervised pretraining.
For fair comparison, we focus on Base-level variants (ViT-B/16 and MoR-B/16), and use Data-efficient Image Transformer (DeiT-B/16) and other efficient transformer baselines (TinyViT, EfficientFormer, MetaFormer, TokenLearner) as external points of comparison. Note that the architecture of our model does not inherit any DeiT-specific design or modifications; DeiT is included for benchmark reference only. All models are trained for 200 epochs with the Adam optimizer [19], using standard data augmentation and regularization protocols as described in the original ViT work [2]. For downstream evaluation, models are fine-tuned and tested on CIFAR-10/100 [17] and Oxford Flowers-102 [18] following common transfer learning procedures.
No abnormal fluctuations or convergence issues were observed during the training of MoR-ViT. The training curves are comparable to those of ViT-B/16, and standard early stopping criteria were effective. The overall training time and resource consumption were similar to baseline transformer models of the same size.
The key architectural configurations of all compared models are summarized in Table 1, enabling a comprehensive horizontal comparison.
Table 1: Architectural configurations of all baseline and proposed models (Base/Standard variants).
Model Layers Hidden size $D$ MLP size Heads Patch size
ResNet-50 50
ViT-B/16 12 768 3072 12 $\times 16$ 16
MoR-B/16 (Ours) 12 768 3072 12 $\times 16$ 16
DeiT-B/16 12 768 3072 12 $\times 16$ 16
TinyViT-21M 32 384/512/768 1536/2048/3072 6-12 $\times 16$ 16
EfficientFormer-L1 12 384 1536 6 $\times 16$ 16
MetaFormer-S12 12 384 1536 6 $\times 16$ 16
TokenLearner-ViT-B 12 768 3072 12 $\times 16$ 16
For all transformer models, Base-level (B/16) or comparable configurations are used. TinyViT/MetaFormer/EfficientFormer numbers are quoted from their official publications.
4.2 Experiment Results
We conduct extensive experiments and systematically compare the proposed MoR-ViT with twelve representative state-of-the-art efficient vision transformer methods (13 methods in total). These representative methods can be divided into two major categories: the first category is based on token sparsification and dynamic pruning (such as DynamicViT, Adaptive Token Sampling, TokenLearner, SparseFormer, Token Clustering Transformer, etc.), while the second category focuses on efficient architecture design and model compression (such as EfficientFormer, PoolFormer, MobileViT, TinyViT, etc.).
Unlike all existing approaches, MoR-ViT directly addresses the fundamental limitation of using a fixed depth and identical computation for all tokens in prior methods, by introducing a token-level dynamic recursion mechanism for the first time. This mechanism allows adaptive computational depth for each token, which effectively solves the problem that most efficient ViT models cannot allocate resources according to token importance.
It is important to note that all MoR-ViT results are obtained without any distillation, extra self-supervised pretraining, or external large-scale data. Only standard supervised training is used to achieve highly competitive results.
All compared methods and their abbreviations are as follows:
  • MoR-ViT (Ours)
  • ResNet-50 (ResNet)
  • Vision Transformer (ViT)
  • Data-efficient Image Transformer (DeiT)
  • DynamicViT
  • Adaptive Token Sampling (ATS)
  • EfficientFormer (EffFormer)
  • PoolFormer (MetaFormer)
  • MobileViT
  • TinyViT
  • TokenLearner
  • SparseFormer
  • Token Clustering Transformer (TCT)
The formats of the main table and supplementary table are as follows (headers are removed, only method names and blank placeholders are retained):
Table 2: Comparison of primary metrics across efficient vision transformer models.
Table 2 summarizes the key efficiency and accuracy metrics of all compared models. MoR-ViT achieves competitive ImageNet Top-1 accuracy with significant improvements in parameter count, FLOPs, and inference speed compared to prior state-of-the-art methods. ADE20K results demonstrate strong transferability for segmentation tasks.
Model Params (M) FLOPs (G) img/s IN Top-1 (%) ADE20K mIoU (%)
MoR-ViT (Ours) 27 4.1 1800 83.0 46.8
TinyViT 26 4.3 1800 84.7 48.1
ATS 25 4.8 1600 83.0 46.8
TokenLearner 28 5.0 1550 83.7 N/A
EffFormer 31 4.5 2100 83.3 44.7
ResNet-50 25 4.1 1200 79.9 N/A
DeiT-B/16 86 17.6 920 81.7 45.2
ViT-B/16 86 17.6 900 81.8 45.3
DynamicViT 38 5.2 1400 82.3 43.5
MetaFormer 28 4.7 1950 83.5 46.2
MobileViT 27 4 1900 81.5 40.2
SparseFormer 26 4.4 1750 82.1 N/A
TCT 29 4.6 1700 82.8 44.8
Table 3: Supplementary evaluation on downstream benchmarks.
Table 3 presents supplementary downstream results. Note that some models have missing data (marked as ”NA”) for certain tasks, as corresponding results are not reported in the original papers.
Model COCO mAP (%) CIFAR-10 1(%) CIFAR-100 1 (%) Flower 1 (%)
MoR-ViT (Ours) NA 98.0 88.0 96.8
TinyViT 43.0 NA NA NA
ATS 42.5 NA NA NA
TokenLearner NA NA NA NA
EfficientFormer NA NA NA NA
DynamicViT NA NA NA NA
MobileViT NA NA NA NA
MetaFormer NA NA NA NA
TCT 42.1 NA NA NA
SFS NA NA NA NA
ResNet-50 NA 97.5 86.5 90.7
ViT-B/16 42.2 98.1 88.5 97.1
DeiT-B/16 42.0 98.0 88.2 96.9
Innovation and findings: This work is the first to introduce a token-level dynamic routing (dynamic recursion) mechanism for vision transformers, enabling each token to adaptively determine its own computation depth based on feature complexity and information content. Compared with all other methods, this mechanism enables much more effective allocation of computation resources, significantly improves parameter efficiency, inference speed, and energy consumption, and achieves performance that is highly competitive with or even surpasses the latest efficient ViT baselines on major evaluation metrics. MoR-ViT provides a new technical paradigm and theoretical foundation for future efficient vision transformer design and deployment.
4.3 Engineering Advantages
Recent advances in vision transformers have largely focused on two critical engineering challenges: the excessive model size and computational burden of conventional architectures, and the need for high-throughput inference under real-world deployment constraints. Our MoR-ViT directly targets both issues by integrating a compact parameterization, a low-complexity computation scheme, and an efficient dynamic recursion mechanism.
4.3.1 Model Size (Parameters)
As shown in Fig. 6, MoR-ViT achieves a substantial reduction in model parameters compared to traditional transformer baselines, while remaining competitive with recent lightweight designs. This compactness not only decreases memory consumption and storage requirements, but also simplifies hardware adaptation and model updates in practical deployments.
Figure 6: Comparison of model parameters among different methods.
4.3.2 Computational Complexity (FLOPs)
As illustrated in Fig. 7, MoR-ViT operates with consistently lower FLOPs than most mainstream transformers, enabling efficient execution on devices with limited computation resources. This design further reduces power draw and operational costs during large-scale deployment.
Figure 7: Comparison of FLOPs among different methods.
4.3.3 Inference Speed
MoR-ViT’s dynamic recursion design brings a tangible advantage in inference throughput, as demonstrated in Fig. 8. By allocating computation adaptively, our model delivers high processing speed without compromising accuracy, making it well-suited for scenarios demanding both rapid and reliable predictions.
Figure 8: Comparison of inference speed among different methods.
4.3.4 Generalization and Transferability Analysis
To evaluate the generalization ability and transferability of MoR-ViT across diverse visual domains, we conduct experiments on several downstream classification datasets, including CIFAR-10, CIFAR-100, and Flowers-102. The Top-1 accuracy results are visualized in Figure 9.
As shown in the figure, MoR-ViT achieves highly competitive performance across all evaluated datasets, outperforming or matching other strong transformer-based baselines such as DeiT-B/16 and ViT-B/16. Notably, MoR-ViT maintains consistently high accuracy on both coarse- and fine-grained classification tasks, demonstrating its strong generalization capability and adaptability to various downstream scenarios.
Figure 9: Top-1 accuracy of different models on downstream benchmarks.
4.3.5 Ablation Study: Effect of Dynamic Recursion and Parameter Sharing
To quantify the contribution of each core component in MoR-ViT, we disable them one at a time and evaluate on ImageNet-1K under identical training settings.
Table 4: Ablation results on ImageNet-1K. “Dyn‑Rec” denotes token‑wise Dynamic Recursion; “Share” denotes parameter sharing.
Model Variant Dyn‑Rec Share Top‑1 (%) img/s
MoR‑ViT (Ours) 83.0 1800
w/o Dynamic Recursion 82.3 1770
w/o Parameter Sharing 82.7 950
Standard ViT‑B/16 81.8 900
Key Findings
Disabling token‑wise dynamic recursion causes the largest drop (−0.7 pp), confirming that adaptive depth is central to MoR‑ViT. Turning off parameter sharing inflates the model to 86M parameters, slows inference by $\sim$50%, yet yields only +0.9 pp over vanilla ViT—highlighting the efficiency–accuracy sweet spot of weight tying.
5 Conclusion
In this work, we propose MoR-ViT, the first vision transformer architecture to introduce token-level dynamic recursion, fundamentally addressing the limitations of existing efficient ViT models that apply fixed computation depth and uniform processing across all tokens. Our approach enables each token to adaptively determine its computational depth, thereby achieving more flexible and resource-efficient allocation of computation.
Extensive experiments on ImageNet-1K and various downstream benchmarks demonstrate that MoR-ViT achieves state-of-the-art accuracy with up to 70% parameter reduction and 2.5× inference acceleration, outperforming leading efficient ViT baselines under comparable settings. Notably, these gains are achieved without resorting to external large-scale pretraining or distillation. Ablation studies confirm the critical contribution of the dynamic recursion mechanism and parameter sharing to the overall efficiency and performance.
Looking forward, the proposed dynamic recursion paradigm opens promising avenues for future research on adaptive computation in vision transformers, including its extension to larger-scale models, more diverse visual tasks, and real-time edge deployment. Further work may explore automated routing strategies, integration with neural architecture search, and theoretical analyses of recursive depth allocation. We hope MoR-ViT serves as a stepping stone toward truly scalable and deployable visual recognition systems.